# Groupoid Graded Semisimple Rings

### v1

@@ 3.3. Graded matrix rings
Now we turn our attention to the representation of homomorphisms with degree of pseudo-free modules as matrices. For that, first we give some technical definitions that will be helpful when defining the graded rings of matrices (not necessarily of finite size) we will consider in later sections.
Let $I$ be a non-empty set and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of nonempty subsets. We say that $\overline{\Sigma}$ is d-finite if, for each $e\in\Gamma_{0}$, the set
@@
    In Proposition 3.7(1), if the sequence $\overline{\Sigma}$ is cf-matricial for $R$, but not matricial for $R$, then $\operatorname{M}_{I}(R)$ is $\Gamma$-graded ring which is not object unital.
@@ 3.4. Graded endomorphism rings as graded matrix rings
Let $I$, $J$ be nonempty sets and $\{M_{j}\colon j\in J\}$, $\{N_{i}\colon i\in I\}$ be families of $\Gamma$-graded $R$-modules. Set $M=\bigoplus\limits_{j\in J}M_{j}$ and $N=\bigoplus\limits_{i\in I}N_{i}$.
@@
Let $I$, $J$ be nonempty sets, $\{M_{j}\colon j\in J\}$ and $\{N_{i}\colon i\in I\}$ be families of $\Gamma$-graded $R$-modules and $M=\bigoplus\limits_{j\in J}M_{j}$, $N=\bigoplus\limits_{i\in I}N_{i}$. Suppose that, for all $j\in J$ and $g\in\operatorname{h}(\operatorname{HOM}(M_{j},N))$, there exists a finite $I_{g}\subseteq I$ such that $\operatorname{im}g\subseteq\bigoplus\limits_{i\in I_{g}}N_{i}$. Consider the natural inclusions $\rho_{j}\colon M_{j}\rightarrow M$, $\rho_{i}^{\prime}\colon N_{i}\rightarrow N$ and the natural projections $\pi_{j}\colon M\rightarrow M_{j}$, $\pi^{\prime}_{i}\colon N\rightarrow N_{i}$ for all $i\in I$, $j\in J$. The following statements hold true.
@@
    The natural map
    $\operatorname{HOM}(M,N)\rightarrow\operatorname{\mathbb{H}}_{I\times J}(M,N),\quad f\mapsto(\pi^{\prime}_{i}f\rho_{j})_{ij},$
    is a gr-isomorphism of $\Gamma$-graded additive groups.
@@
    The natural map
    $\operatorname{END}(M)\rightarrow\operatorname{\mathbb{H}}_{J\times J}(M,M),\quad f\mapsto(\pi_{i}f\rho_{j})_{ij},$
    is a gr-isomorphism of $\Gamma$-graded rings.
@@
First we prove that $\Phi$ is well defined. For that it is enough to show that $\Phi$ is well defined for homogeneous elements. Let $\gamma\in\Gamma$ and $f\in\operatorname{HOM}_{R}(M,N)_{\gamma}$. For each $i\in I,j\in J$ we have that $\pi^{\prime}_{i}f\rho_{j}\in\operatorname{HOM}_{R}(M_{j},N_{i})_{\gamma}$. Fix now $j\in J$. By hypothesis, there exists a finite $I_{f\rho_{j}}\subseteq I$ such that $\operatorname{im}(f\rho_{j})\subseteq\bigoplus\limits_{i\in I_{f\rho_{j}}}N_{i}$. Hence, $\pi^{\prime}_{i}f\rho_{j}\neq 0$ implies $i\in I_{f\rho_{j}}$ and it follows that the sequence $(\pi^{\prime}_{i}f\rho_{j})_{i\in I}$ is almost zero.
@@
Let $I$ be a nonempty set and $\{M_{i}\colon i\in I\}$ be a family of $\Gamma$-graded $R$-modules such that $\operatorname{HOM}(M_{j},M_{i})=0$ for different $i,j\in I$. Consider $M=\bigoplus\limits_{i\in I}M_{i}$. Then
@@
Let $M$ be a $\Gamma$-graded $R$-module and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of nonempty subsets of $\Gamma$. We will denote
@@
(4) follows from (2) and Lemma 3.4. ∎
@@
Let $I$ be a nonempty set, $e\in\Gamma_{0}$ and $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e\Gamma)^{I}$. The following statements hold true.
@@
    Suppose $\operatorname{supp}(R)\subseteq e\Gamma e$. If $\overline{\sigma}$ is $d$-finite, then $\overline{\sigma}$ is fully matricial for $R$ and therefore
    $\operatorname{END}(R(\overline{\sigma}))\cong_{gr}\operatorname{CFM}_{I}(R)(\overline{\sigma})=\operatorname{M}_{I}(R)(\overline{\sigma})=\operatorname{M}_{I}(R),$
    where the last equality is as rings.
@@
    If $\Sigma_{i}$ is $d$-unique and $r$-unique for $R$ for all $i=1,\dotsc,n$, then $\overline{\Sigma}$ is matricial for $R$ and
    $\operatorname{END}(R(\overline{\Sigma}))\cong_{gr}\operatorname{CFM}_{n}(R)(\Sigma_{1},\dotsc,\Sigma_{n})=\operatorname{M}_{n}(R)(\Sigma_{1},\dotsc,\Sigma_{n}).$
    If, moreover, $\Sigma_{i}$ is fully $r$-unique for $R$ for all $i=1,\dotsc,n$, then we also have the equality as rings $\operatorname{M}_{n}(R)=\operatorname{M}_{n}(R)(\Sigma_{1},\dotsc,\Sigma_{n})$.
@@ 3.5. Some results concerning the opposite ring
  In this subsection we prove a result concerning the opposite ring. It will be helpful in order to obtain results about left $R$-modules from the ones obtained for right $R$-modules.
@@
If $g\colon L\rightarrow L^{\prime}$ is homomorphism of left $R$-modules, we define $\hat{g}\colon L^{op}\rightarrow(L^{\prime})^{op}$ by $\hat{g}(x)=(x)g$ for all $x\in L$. Since
@@
Let $L$ be a $\Gamma$-graded left $R$-module and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of $d$-unique sets. The following assertions hold true.
@@
    If $\overline{\Sigma}$ is matricial for $R$, then $\overline{\Sigma}$ is matricial for $R^{op}$ and
    $\operatorname{M}_{I}(R)(\overline{\Sigma})^{op}\cong_{gr}\operatorname{M}_{I}(R^{op})(\overline{\Sigma})$
    via the homomorphism defined by transposition of matrices.
@@
(1) The first affirmation it follows from $\Gamma^{\prime}_{0}(R)=\Gamma^{\prime}_{0}(R^{op})$. Let
@@
Hence, $A^{t}\in\operatorname{M}_{I}(R^{op})(\overline{\Sigma})_{\gamma}$. Therefore $\Phi$ is well defined and it is also a gr-homomorphism of additive groups. Given $A,B\in\operatorname{M}_{I}(R)(\overline{\Sigma})^{op}$, we have
$\Phi(A\cdot^{op}B)=\Phi(BA)=(BA)^{t}\stackrel{{\scriptstyle(*)}}{{=}}A^{t}B^{t}=\Phi(A)\Phi(B),$
where, in ()(*)( ∗ ), we have used that the $(i,j)$-entry of $(BA)^{t}$ is
@@
It implies that $(\operatorname{END}_{R}L)^{op}\to\operatorname{END}_{R^{op}}(L^{op})$, $g\mapsto\hat{g}$, is well defined and it is bijective. Finally, observe that if $g_{1},g_{2}\in\operatorname{END}_{R}(L)$, then $\widehat{g_{2}\circ^{op}g_{1}}=\widehat{g_{1}g_{2}}=\hat{g}_{2}\hat{g}_{1}$, as desired.
@@ 3.6. Some isomorphisms between categories of graded modules
Let $R$ be a $\Gamma$-graded ring. The category whose objects are the $\Gamma$-graded right (resp. left) $R$-modules and morphisms are gr-homomorphisms will be denoted by $\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma-R\operatorname{\!\textrm{-}gr}$). If $e\in\Gamma_{0}$ is fixed, then the full subcategory of $\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma-R\operatorname{\!\textrm{-}gr}$) whose objects are the $\Gamma$-graded right (resp. left) $R$-modules $M$ such that $\operatorname{supp}M\subseteq e\Gamma$ (resp. $\operatorname{supp}M\subseteq\Gamma e$) will be denoted by $e\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma e-R\operatorname{\!\textrm{-}gr}$). We denote by $-\operatorname{\varepsilon\Gamma}-\operatorname{mod\textrm{-}\!}$-$R$ (resp. $\operatorname{\Gamma\varepsilon}-R\operatorname{\!\textrm{-}mod}$) the category whose objects are the $\Gamma$-graded right (resp. left) $R$-modules $M$ for which there exists $\varepsilon_{M}\in\Gamma_{0}$ satisfying $\operatorname{supp}M\subseteq\varepsilon_{M}\Gamma$ (resp. $\operatorname{supp}M\subseteq\Gamma\varepsilon_{M}$) and, for objects $M$ and $N$, the set of morphisms from $M$ to $N$ is $\operatorname{HOM}_{R}(M,N)$. Note that if $M=M(e)$ for some $e\in\Gamma_{0}$, then $\mathds{1}_{e}$ is the unity of the ring $\operatorname{HOM}_{R}(M,M)$. In some cases, we will refer to objects of $e\Gamma-\operatorname{gr\textrm{-}\!}R$ as $e\Gamma$-graded $R$-modules.
We want to describe categories of left modules using categories of right modules. Let $R$ be a $\Gamma$-graded ring. In this subsection we will regard the ring $R^{op}$ as a $\Gamma^{op}$-graded ring via $(R^{op})_{\gamma^{op}}:=R_{\gamma}$, where $\Gamma^{op}$ is the opposite category of $\Gamma$. In this context, if $L$ is a $\Gamma$-graded left $R$-module, then $L^{op}$ will be regarded as a $\Gamma^{op}$-graded right $R^{op}$-module via $(L^{op})_{\gamma^{op}}:=L_{\gamma}$. Note that
$(L^{op})_{\gamma^{op}}\cdot^{op}(R^{op})_{\delta^{op}}=R_{\delta}L_{\gamma}\subseteq L_{\delta\gamma}=(L^{op})_{(\delta\gamma)^{op}}=(L^{op})_{\gamma^{op}\delta^{op}}.$
@@
Let $R$ be a $\Gamma$-graded ring. We have the following isomorphisms between categories.
@@
    $\Gamma-R\operatorname{\!\textrm{-}gr}\cong\Gamma^{op}-\operatorname{gr\textrm{-}\!}R^{op}$.
@@
    $\Gamma e-R\operatorname{\!\textrm{-}gr}\cong e^{op}\Gamma^{op}-\operatorname{gr\textrm{-}\!}R^{op}$.
@@
    $-\operatorname{\Gamma\varepsilon}-R\operatorname{\!\textrm{-}mod}\cong\operatorname{\varepsilon\Gamma}^{op}-\operatorname{mod\textrm{-}\!}$-$R^{op}$.
@@
In all three cases, we define a functor from the corresponding category of graded left $R$-modules to the category of graded right $R$-modules. This functor is defined on the objects by $L\mapsto L^{op}$, and sends any morphism $g:L\to L^{\prime}$ to the morphism $\hat{g}:L^{op}\rightarrow(L^{\prime})^{op}$ defined on Section 3.5. In the third case, observe that if $g$ is of degree $\gamma$, then $\hat{g}$ is of degree $\gamma^{op}$. ∎

### v2

@@ 3.3. Graded matrix rings
Now we turn our attention to the representation of homomorphisms with degree of pseudo-free modules as matrices. For this purpose, given a ring $R$ and a set $I$, we consider two rings of $I\times I$ matrices with entries in $R$. The first one is $\operatorname{CFM}_{I}(R)$, consisting of the column finite $I\times I$ matrices with entries in $R$. The second one is $\operatorname{M}_{I}(R)$, consisting of the $I\times I$ matrices with only a finite number of nonzero entries in $R$. In the ungraded context, if $R$ is a unital ring these matrix rings can be identified with certain rings of endomorphisms of free $R$-modules, as the next example shows.
Example 3.6.
Let $R$ be a unital ring and let $M$ be a free right $R$-module with basis $\mathcal{B}=\{e_{i}\colon i\in I\}.$ Let $f\in\operatorname{End}_{R}(M)$. We can associate to $f$ the matrix $\left[f\right]_{\mathcal{B}}\in\operatorname{CFM}_{I}(R)$, whose $i$-th column consists of the coordenates of $f(e_{i})$ with respect to the basis $\mathcal{B}$ for each $i\in I$. Then the map $\operatorname{End}_{R}(M)\rightarrow\operatorname{CFM}_{I}(R)$, given by $f\mapsto\left[f\right]_{\mathcal{B}}$, defines a ring isomorphism. Under this isomorphism, the subring $\operatorname{M}_{I}(R)$ of $\operatorname{CFM}_{I}(R)$ is then identified with the subring of $\operatorname{End}_{R}(M)$ consisting of the endomorphisms $f$ such that the set $I_{f}=\{i\in I\colon f(e_{i})\neq 0\}$ is finite. Notice that if $I$ is finite, then $\operatorname{CFM}_{I}(R)=\operatorname{M}_{I}(R)$. ∎
In this subsection, we aim to generalize Example 3.6 to the groupoid graded context. To this end, we have to define first the graded objects that will play the role of $\operatorname{CFM}_{I}(R)$ and $\operatorname{M}_{I}(R)$.
Let $I$ be a non-empty set and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of non-empty subsets. We say that $\overline{\Sigma}$ is d-finite if, for each $e\in\Gamma_{0}$, the set
@@
    In Proposition 3.8(1), if the sequence $\overline{\Sigma}$ is cf-matricial for $R$, but not matricial for $R$, then $\operatorname{M}_{I}(R)$ is a $\Gamma$-graded ring which is not object unital.
@@ 3.4. Graded endomorphism rings as graded matrix rings
Let $I$, $J$ be non-empty sets and $\{M_{j}\colon j\in J\}$, $\{N_{i}\colon i\in I\}$ be families of $\Gamma$-graded $R$-modules. Set $M=\bigoplus\limits_{j\in J}M_{j}$ and $N=\bigoplus\limits_{i\in I}N_{i}$.
@@
Let $I$, $J$ be non-empty sets, $\{M_{j}\colon j\in J\}$ and $\{N_{i}\colon i\in I\}$ be families of $\Gamma$-graded $R$-modules and $M=\bigoplus\limits_{j\in J}M_{j}$, $N=\bigoplus\limits_{i\in I}N_{i}$. Suppose that, for all $j\in J$ and $g\in\operatorname{h}(\operatorname{HOM}(M_{j},N))$, there exists a finite $I_{g}\subseteq I$ such that $\operatorname{im}g\subseteq\bigoplus\limits_{i\in I_{g}}N_{i}$. Consider the natural inclusions $\rho_{j}\colon M_{j}\rightarrow M$, $\rho_{i}^{\prime}\colon N_{i}\rightarrow N$ and the natural projections $\pi_{j}\colon M\rightarrow M_{j}$, $\pi^{\prime}_{i}\colon N\rightarrow N_{i}$ for all $i\in I$, $j\in J$. The following statements hold true.
@@
    The natural map HOM(M,N)H_I×J(M,N),  f(π’_ifρ_j)_ij, is a gr-isomorphism of $\Gamma$-graded additive groups.
@@
    The natural map END(M)H_J×J(M,M),  f(π_ifρ_j)_ij, is a gr-isomorphism of $\Gamma$-graded rings.
@@
First we prove that $\Phi$ is well-defined. For that it is enough to show that $\Phi$ is well defined for homogeneous elements. Let $\gamma\in\Gamma$ and $f\in\operatorname{HOM}_{R}(M,N)_{\gamma}$. For each $i\in I,j\in J$ we have that $\pi^{\prime}_{i}f\rho_{j}\in\operatorname{HOM}_{R}(M_{j},N_{i})_{\gamma}$. Fix now $j\in J$. By hypothesis, there exists a finite $I_{f\rho_{j}}\subseteq I$ such that $\operatorname{im}(f\rho_{j})\subseteq\bigoplus\limits_{i\in I_{f\rho_{j}}}N_{i}$. Hence, $\pi^{\prime}_{i}f\rho_{j}\neq 0$ implies $i\in I_{f\rho_{j}}$ and it follows that the sequence $(\pi^{\prime}_{i}f\rho_{j})_{i\in I}$ is almost zero.
@@
Let $I$ be a non-empty set and $\{M_{i}\colon i\in I\}$ be a family of $\Gamma$-graded $R$-modules such that $\operatorname{HOM}(M_{j},M_{i})=0$ for different $i,j\in I$. Consider $M=\bigoplus\limits_{i\in I}M_{i}$. Then
@@
Let $M$ be a $\Gamma$-graded $R$-module and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of non-empty subsets of $\Gamma$. We will denote
@@
(4) By Lemma 3.4, $R\cong_{gr}\operatorname{END}(R_{R})$. Thus, if we make $M=R_{R}$ in (2), we obtain $\operatorname{END}(R(\overline{\Sigma}))\cong_{gr}\operatorname{M}_{I}(\operatorname{END}(R_{R}))(\overline{\Sigma})\cong_{gr}\operatorname{M}_{I}(R)(\overline{\Sigma})$ and the first part of the statement is proved.
Suppose now that $\overline{\Sigma}$ is fully matricial for $R$. Again, if we make $M=R_{R}$ in (2), we get $\operatorname{END}(R(\overline{\Sigma}))$ is isomorphic to $\operatorname{M}_{I}(\operatorname{END}(R_{R}))$ as graded rings (with the induced grading from $\operatorname{M}_{I}(\operatorname{END}(R_{R}))(\overline{\Sigma})$). Therefore, $\operatorname{END}(R(\overline{\Sigma}))$ is isomorphic to $\operatorname{M}_{I}(R)$ as graded rings (with the induced grading from $\operatorname{M}_{I}(R)(\overline{\Sigma})$).∎
@@
Let $I$ be a non-empty set, $e\in\Gamma_{0}$ and $\overline{\sigma}=(\sigma_{i})_{i\in I}\in(e\Gamma)^{I}$. The following statements hold true.
@@
    Suppose $\operatorname{supp}(R)\subseteq e\Gamma e$. If $\overline{\sigma}$ is $d$-finite, then $\overline{\sigma}$ is fully matricial for $R$ and therefore END(R(σ))_gr CFM_I(R)(σ)= M_I(R)(σ)=M_I(R), where the last equality is as rings.
@@
    If $\Sigma_{i}$ is $d$-unique and $r$-unique for $R$ for all $i=1,\dotsc,n$, then $\overline{\Sigma}$ is matricial for $R$ and END(R(Σ))_gr CFM_n(R)(Σ_1,,Σ_n)=M_n(R)(Σ_1,,Σ_n). If, moreover, $\Sigma_{i}$ is fully $r$-unique for $R$ for all $i=1,\dotsc,n$, then we also have the equality as rings $\operatorname{M}_{n}(R)=\operatorname{M}_{n}(R)(\Sigma_{1},\dotsc,\Sigma_{n})$.
@@ 3.5. Some results concerning the opposite ring
The aim of this section is to present techniques for deriving results about gr-semisimple left $R$-modules from those obtained for gr-semisimple right $R$-modules.
@@
Let $g\colon L\rightarrow L^{\prime}$ be a homomorphism of left $R$-modules. Recall that, since $L$ is a left $R$-module, the homomorphism $g$ acts on elements of $L$ from the right by convention. We define $\hat{g}\colon L^{op}\rightarrow(L^{\prime})^{op}$ by $\hat{g}(x)=(x)g$ for all $x\in L$. Since
@@
Let $L$ be a $\Gamma$-graded left $R$-module and $\overline{\Sigma}=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a sequence of $d$-unique sets. The following assertions hold:
@@
    If $\overline{\Sigma}$ is matricial for $R$, then $\overline{\Sigma}$ is matricial for $R^{op}$ and M_I(R)(Σ)^op_grM_I(R^op)(Σ) via the homomorphism defined by transposition of matrices.
@@
(1) The first part follows from $\Gamma^{\prime}_{0}(R)=\Gamma^{\prime}_{0}(R^{op})$. Let
@@
Hence, $A^{t}\in\operatorname{M}_{I}(R^{op})(\overline{\Sigma})_{\gamma}$. Therefore, $\Phi$ is well-defined and it is also a gr-homomorphism of additive groups. Given $A,B\in\operatorname{M}_{I}(R)(\overline{\Sigma})^{op}$, we have
$\Phi(A\cdot^{op}B)=\Phi(BA)=(BA)^{t}\mathrel{{\mathop{=}\limits^{(*)}}}A^{t}B^{t}=\Phi(A)\Phi(B),$
where, in ()(*), we have used that the $(i,j)$-entry of $(BA)^{t}$ is
@@
This implies that the map $(\operatorname{END}_{R}L)^{op}\to\operatorname{END}_{R^{op}}(L^{op})$, $g\mapsto\hat{g}$, is well-defined and it is bijective. Finally, observe that if $g_{1},g_{2}\in\operatorname{END}_{R}(L)$, then $\widehat{g_{2}\circ^{op}g_{1}}=\widehat{g_{1}g_{2}}=\hat{g}_{2}\hat{g}_{1}$, as desired.
@@ 3.6. Some isomorphisms between categories of graded modules
Let $R$ be a $\Gamma$-graded ring. The category whose objects are the $\Gamma$-graded right (resp. left) $R$-modules and morphisms are gr-homomorphisms will be denoted by $\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma-R\operatorname{\!\textrm{-}gr}$). If $e\in\Gamma_{0}$ is fixed, then the full subcategory of $\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma-R\operatorname{\!\textrm{-}gr}$) whose objects are the $\Gamma$-graded right (resp. left) $R$-modules $M$ such that $\operatorname{supp}M\subseteq e\Gamma$ (resp. $\operatorname{supp}M\subseteq\Gamma e$) will be denoted by $e\Gamma-\operatorname{gr\textrm{-}\!}R$ (resp. $\Gamma e-R\operatorname{\!\textrm{-}gr}$). We denote by $\operatorname{\varepsilon\Gamma}-\operatorname{mod\textrm{-}\!}$-$R$ (resp. $\operatorname{\Gamma\varepsilon}-R\operatorname{\!\textrm{-}mod}$) the category whose objects are the $\Gamma$-graded right (resp. left) $R$-modules $M$ for which there exists $\varepsilon_{M}\in\Gamma_{0}$ satisfying $\operatorname{supp}M\subseteq\varepsilon_{M}\Gamma$ (resp. $\operatorname{supp}M\subseteq\Gamma\varepsilon_{M}$) and, for objects $M$ and $N$, the set of morphisms from $M$ to $N$ is $\operatorname{HOM}_{R}(M,N)$. Note that if $M=M(e)$ for some $e\in\Gamma_{0}$, then $\mathds{1}_{e}$ is the unity of the ring $\operatorname{HOM}_{R}(M,M)$. In some cases, we will refer to objects of $e\Gamma-\operatorname{gr\textrm{-}\!}R$ as $e\Gamma$-graded $R$-modules.
We aim to describe categories of left modules using categories of right modules. One way to proceed is to induce gradings on opposite rings and modules using $\Gamma$ as in the previous subsection. An alternative approach, which we adopt here, is to work with $\Gamma^{op}$, the opposite category of $\Gamma$, which is again a groupoid and is called the opposite groupoid of $\Gamma$. It comes equipped with a natural bijection $\Gamma\rightarrow\Gamma^{op}$, given by $\gamma\mapsto\gamma^{o}$, where $d(\gamma^{o})=r(\gamma)^{o}$ and $r(\gamma^{o})=d(\gamma)^{o}$. Furthermore, if $\gamma,\delta\in\Gamma$ are such that $\gamma\delta$ is defined in $\Gamma$, then $(\gamma\delta)^{o}=\delta^{o}\gamma^{o}$ and $(\gamma^{-1})^{o}=(\gamma^{o})^{-1}$.
Let $R$ be a $\Gamma$-graded ring. The next result considers the ring $R^{op}$ as a $\Gamma^{op}$-graded ring via the grading $(R^{op})_{\gamma^{o}}:=R_{\gamma}$. In this context, if $L$ is a $\Gamma$-graded left $R$-module, then $L^{op}$ will be regarded as a $\Gamma^{op}$-graded right $R^{op}$-module via $(L^{op})_{\gamma^{o}}:=L_{\gamma}$. Note that
$(L^{op})_{\gamma^{o}}\cdot^{op}(R^{op})_{\delta^{o}}=R_{\delta}L_{\gamma}\subseteq L_{\delta\gamma}=(L^{op})_{(\delta\gamma)^{o}}=(L^{op})_{\gamma^{o}\delta^{o}}.$
@@
Let $R$ be a $\Gamma$-graded ring. We have the following category isomorphisms:
@@
    $\Gamma-R\operatorname{\!\textrm{-}gr}\cong\Gamma-\operatorname{gr\textrm{-}\!}R^{op}\cong\Gamma^{op}-\operatorname{gr\textrm{-}\!}R^{op}$.
@@
    $\Gamma e-R\operatorname{\!\textrm{-}gr}\cong e\Gamma-\operatorname{gr\textrm{-}\!}R^{op}\cong e^{o}\Gamma^{op}-\operatorname{gr\textrm{-}\!}R^{op}$.
@@
    $\operatorname{\Gamma\varepsilon}-R\operatorname{\!\textrm{-}mod}\cong\operatorname{\varepsilon\Gamma}-\operatorname{mod\textrm{-}\!}$-$R^{op}\cong\operatorname{\varepsilon\Gamma}^{op}-\operatorname{mod\textrm{-}\!}$-$R^{op}$.
@@
In all three cases, we define a functor from the corresponding category of $\Gamma$-graded left $R$-modules to the category of $\Gamma$-graded (or $\Gamma^{op}$-graded) right $R$-modules. This functor is defined on objects by $L\mapsto L^{op}$, and sends any morphism $g:L\to L^{\prime}$ to the morphism $\hat{g}:L^{op}\rightarrow(L^{\prime})^{op}$ defined on Section 3.5. In the third case, observe that if $g$ is of degree $\gamma$ in $\operatorname{\Gamma\varepsilon}-R\operatorname{\!\textrm{-}mod}$, then $\hat{g}$ is of degree $\gamma^{o}$ in $\operatorname{\varepsilon\Gamma}^{op}-\operatorname{mod\textrm{-}\!}$-$R^{op}$ and of degree $\gamma^{-1}$ in $\operatorname{\varepsilon\Gamma}-\operatorname{mod\textrm{-}\!}$-$R^{op}$. ∎